OpenStudy (anonymous):

solve the equation. Check the solution. -2/x+4=4/x+3

OpenStudy (anonymous):

(-2/x)+4=(4/x)+3 Am i correct ?

OpenStudy (anonymous):

4-3=4/x+2/x 1=6/x x=6